OpenStudy (anonymous):

help please,fan and medal will be awarded to the best/helpful answer! :3 the density of a gasoline at room temperate is 0.70g/mL. if the gas tank of a car holds 12.0 gallons of gasoline, what is the mass of a full tank of gasoline? note that 1 gallon = 3.7854L

OpenStudy (anonymous):

volume of tank = 12 * 3,7845 density = m/V m = ?!

OpenStudy (anonymous):

so since m = density x volume, do i just multiply (0.070g/mL) (12*3,7854L)?

OpenStudy (anonymous):

check your units!!! are all units in g or kg, are all units in mL or L?

OpenStudy (anonymous):

g/mL. L?

OpenStudy (anonymous):

convert the volume to ml, that should be easiest way...

OpenStudy (anonymous):

0.00007g/L

OpenStudy (anonymous):

1 gallon = 3.7854L 12 gallons = 45.4248 L density = 0,7 g/ml = 0,7 kg/L density = m/V m= density * V

OpenStudy (anonymous):

oh wait. my bad loool, density should be 70g/L

OpenStudy (anonymous):

ok then just recalculate it... but i gave you entire solution you just need to put in numbers and mind units...

OpenStudy (anonymous):

thanks! i was having a hard time grasping the concept so i got my teacher to help me today :)